A concave mirror mirror of focal length 20cm is cut into two parts from the middle and the two parts are moved perpendicularly by a distance 1 mm from the previous principal axis AB. Find the distance between the images formed by the two parts.

Grade:11

1 Answers

Vikas TU
6 years ago
Here central length will be same as before that is 20 cm = 0.2 m 
 
The question separation will be = 0.01 m 
 
At that point from the mirror equation 
 
1/v+1/u = 1/f 
 
1/v-100 = 5 
 
1/v = 105 
 
v = 1/105 m